OpenStudy (anonymous):

Simplify (2 {sqrt 5}+ 3 {sqrt 7})^2. Show your work. Justify each step. Thanks in advance for the help!!!

OpenStudy (ash2326):

Is this your question ? \[(2\sqrt 5 + 3\sqrt 7)^2\]

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

Let's use the idenitity, \[(a+b)^2=a^2+2ab+b^2\] here \[a=2\sqrt 5\] \[b=3\sqrt 7\] Can you try ?

OpenStudy (anonymous):

How do I plug that into the equation? I'm very confused....

OpenStudy (ash2326):

\[(2\sqrt 5+3 \sqrt 7)^2=\] \[(2\sqrt 5)^2+2\times(2\sqrt 5)\times(3\sqrt 7)+(3\sqrt 7)^2\] Can you try now?

OpenStudy (anonymous):

Thank you very much for you help! Do I multiply the 2by both 2 and 5 (the part after the first plus sign)? I think I understand how to do the rest of it.

OpenStudy (ash2326):

Only one \[2\times 2\times \sqrt5 \times 3\times \sqrt 7=2\times 2\times 3 \times \sqrt 5\times \sqrt 7 \] \[=12\times \sqrt 5\times \sqrt 7\] There you go \(\uparrow\)

OpenStudy (anonymous):

Thank you!! You've been very helpful :)

OpenStudy (ash2326):

Welcome :)